OpenStudy (anonymous):

x+y=-6, -2x+y=0 solve using Elimination Method

OpenStudy (ajprincess):

x+y=-6-(1) -2x+y=0-(2) subtract (2) from (1). when u do so y will be eliminated and u will get an equation in terms of only x. solve for x. then substitute the value of x in x+y=-6 to find y..

OpenStudy (anonymous):

Thank you so much!!! I am just a little confused on how you got (2) (1) for the problems

OpenStudy (ajprincess):

Oh i jst named the equations.

OpenStudy (anonymous):

hhahahah that confused me!!! thanks for clearning that up :)

OpenStudy (ajprincess):

yw:). Can u do it?

OpenStudy (ajprincess):

Also sorry for the confusion caused.

OpenStudy (anonymous):

yes! im working on it now... if i have problems il let you know!! Iam new to this site and this is awesome!! i WORK a lot and now can get help online while i multi taks at work

OpenStudy (ajprincess):

sure nd welcome to open study.

OpenStudy (anonymous):

I just got the answer and it lookd great!

OpenStudy (ajprincess):

glad to hear that:)